# GS-ROR: 3D Gaussian Splatting for Reflective Object Relighting via SDF Priors

### v2

@@ 7. Conclusion
In this paper, we present a novel framework for real-time reflective object inverse rendering. We design an SDF-aided Gaussian splatting framework, using the mutual supervision of the depth and normal between deferred Gaussians and SDF to improve the quality of geometry from Gaussians. Besides, we propose an SDF-aware pruning strategy with an automatically adjusted threshold, regularizing the position of Gaussian and avoiding the floater artifact. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods without losing efficiency. It is competitive with the NeRF-based methods in terms of quality, with much less training and rendering time. In future work, extending our framework to complex scenarios with multiples objects is a potential direction.

### v3

@@ 7. Conclusion
In this paper, we present a novel framework for real-time reflective object inverse rendering. We design an SDF-aided Gaussian splatting framework, using the mutual supervision of the depth and normal between deferred Gaussians and SDF to improve the geometry quality from Gaussians. Besides, we propose an SDF-aware pruning strategy with an automatically adjusted threshold, regularizing the position of Gaussian and avoiding the floater artifact. Finally, we design a GS-guided SDF enhancement to extract high-quality mesh by finetuning the TensoSDF. Consequently, our method outperforms the existing Gaussian-based inverse rendering methods without losing efficiency. It is competitive with the NeRF-based methods in terms of relighting and mesh quality, with much less training and rendering time. In future work, extending our framework to complex scenarios with multiple objects is a potential direction.